# M. Belianovich, G. E. Fasshauer, A. Narayan, and V. Shankar

### v2

@@ 1 Introduction
[…]
@@
Note: A natural question that arises when reading this work might be “why not use PHS kernels in conjunction with polynomials as global interpolants instead?”. This is indeed a valid choice, but the PHS kernels induce a dense kernel Gramian $A$ in $\mathcal{A}$. Further, due their lack of positive-definiteness, they do not admit efficient linear algebra. Even if one worked with a reduced Schur complement system, our approach is asymptotically faster. Our approach also carries over to any positive-definite kernel, but is particularly efficient for compactly-supported ones. Finally, the storage requirements of the unified interpolant presented in this work are significantly lower than those of PHS kernels with polynomials. We do believe the PHS kernels with polynomials are still an appropriate choice for local interpolation and the generation of RBF-based finite difference (RBF-FD) weights.
@@ 2 Unified interpolation with sparse kernels and polynomials
In the following discussion, let $f:\mathbb{R}^{d}\to\mathbb{R}$ be some target function, and let $X=\{\bm{x}_{k}\}_{k=1}^{N}\subset\mathbb{R}^{d}$ be the set of data sites for interpolation where samples of $f$ are provided. Let $\bm{y}=\left.f\right|_{X}$, i.e., $\bm{y}$ is the vector formed by evaluating $f$ at the $N$ data sites $\bm{x}_{k}$. Further, let $q$ be the separation distance in $X$, and $w$ be the largest pairwise distance. Finally, let $r=1/\varepsilon$ be the support of the Wendland kernel. We are primarily interested in two regimes: (1) $r<q$, where the polynomials dominate the approximation, and (2) $q<r<w$, where both kernels and polynomials contribute to the approximation, but the kernel still produce sparse Gramians. Both these regimes admit fast and specialized numerical linear algebra, which carries over also to the case where $r=w$; we do present some experiments for this latter regime in Section 3.
Note: In $\mathbb{R}^{d}$, we use tensor-products of univariate Legendre polynomials restricted to a total degree index set so that they form a basis for the space of polynomials of total degree $\ell$ in dimension $d$. These are evaluated using the usual three-term recurrence relations. Though this is not necessary for the experiments in this paper (which are all on the unit interval or the unit ball), we rescale all points to the unit interval $[-1,1]^{d}$ so that we can use the standard definitions of Legendre polynomials. We only do this rescaling for the polynomial part of the approximation; the kernels are evaluated on the original point set. We also note that any basis for polynomials can be used instead (even monomials), but we found that tensor products of orthogonal polynomials were easier to work with from a software engineering standpoint.
@@ 2.1.1 Interpolation
When the support $r$ of any Wendland kernel is made smaller than $q$, the shifts of the Wendland kernel each take on the value of $1$ at their center and 00 elsewhere, resembling smooth bump functions. This diagonalizes the kernel Gramian to the $N\times N$ identity matrix $I$; we alternatively refer to the polynomial limit as the diagonal limit for this reason. Thus, the interpolation constraints (2)–(3) enforced at the set of nodes $X$ lead to the following linear system:
@@ 2.1.2 Linear Algebra
    Compute $P=QR$, the (reduced) QR decomposition of the polynomial least-squares matrix $P$ [$O(NM^{2})$].
@@
    Solve the Schur complement system for the polynomial coefficients $\mathbf{d}$ as $\mathbf{d}=R^{-1}Q^{T}\bm{y}$ [$O(N^{2})+O(NM)$].
@@ 2.2.2 Linear Algebra
    Compute $A=LL^{T}$, the Cholesky decomposition of the RBF matrix $A$ [$O(N^{3})]$.
@@
    Compute the matrix $B=L^{-1}P$ and the vector $\mathbf{g}=L^{-1}\bm{y}$ [$O(MN^{2})+O(N^{2})$].
@@
    Compute $B=\tilde{Q}\tilde{R}$, the QR decomposition of $B$ [$O(NM^{2})$].
@@
    Solve for the polynomial coefficients $\mathbf{d}$ as $\mathbf{d}=\tilde{R}^{-1}\tilde{Q}^{T}\mathbf{g}$ [$O(NM)+O(M^{2})$].
@@
    Solve for the RBF coefficients $\bm{c}$ as $\bm{c}=L^{-T}\left(\mathbf{g}-B\mathbf{d}\right)$ [$O(N^{2})$].
@@ 3 Numerical Results
We now present results for interpolation of target functions on Euclidean domains for $d=1,2,3$ and for manifolds $\mathbb{M}\subset\mathbb{R}^{3}$ of co-dimension 1. For $d=1$, the domain is the unit interval $[-1,1]$ and we use Chebyshev extrema (Chebshev-Lobatto nodes) as our interpolation data sites (unless otherwise mentioned). For $d=2$, the domain is the unit disk, and for $d=3$ the unit ball. For both $d=2$ and $d=3$, we use boundary-clustered Poisson disk samples generated using the node generation algorithms from [29] as the interpolation data sites, with the points near the boundary being at $0.75h$ instead of $h$ (where $h$ is the fill distance); two representative node sets are shown in Figure 1. These points are clearly non-uniform. The manifolds are point cloud discretizations of the unit sphere and a torus; we present more details in Section 3.4.
Shape parameter $\epsilon$
For all Wendland kernels, we experimented with two different strategies to control the kernel shape parameter $\varepsilon$:
@@
In order to solve for a shape parameter that meets a target condition number, we use the approach outlined in [27]: we rootfind on $f(\epsilon)=\log 10\left(cond(A(\epsilon)\right)-\log 10(K_{t})$. While other approaches for directly selecting $\epsilon$ based directly on internodal distances may be more applicable in practice, this approach allowed us to automate shape parameter calculations without much trial and error.
Polynomial degree $\ell$
@@
In all experiments, we report the relative $\ell_{2}$ error as $\frac{\|\left.s\right|_{X_{e}}-\left.f\right|_{X_{e}}\|_{2}}{\|\left.f\right|_{X_{e}}\|_{2}}$. For $d=1$, we used $2^{14}$ equispaced points on $[-1,1]$ as the set $X_{e}$. For $d=2,3$, we used densely sampled, quasi-uniform Poisson disk samples with $N_{e}=21748$ and $N_{e}=27987$ respectively, again computed using the node generator from [29]. On manifolds, we used different node generators based on the manifold itself, but set the number of evaluation points to $N_{e}=15000$.
3.1 Univariate functions on $[-1,1]$
@@
We investigate the performance of the hybrid RBF-polynomial interpolation with Wendland kernels of smoothness $C^{2}(\mathbb{R}^{3})$, $C^{4}(\mathbb{R}^{3})$, and $C^{6}(\mathbb{R}^{3})$ for the piecewise-smooth function $f(x)=|x|$). Since this target lacks smoothness, we expect slow convergence rates for all our methods. The increasing polynomial degrees for this case are as follows: $4$, $8$, $16$, $32$, $64$, $128$, $256$.
Figure 2: Relative $\ell_{2}$ error vs. $N^{1/d}$ (here, $N$) for $f(x)=|x|$ using the $C^{2}(\mathbb{R}^{3})$ Wendland kernel. (Left) Fixed-shape (FS) strategy: Shape parameters $\epsilon$ are tuned on the finest grid and reused. (Right) Fixed condition number (FC) strategy: $\epsilon$ is adjusted per grid to maintain fixed condition numbers ($K_{t}=10^{12},10^{8},10^{4}$).
Figure 2 shows that for both FS and FC the relative $\ell_{2}$ error decays algebraically for the $C^{2}(\mathbb{R}^{3})$ kernel; we observed similar results for kernels of higher smoothness as well. This is consistent with the lack of smoothness of the target function. Note that all curves for $K_{t}=10^{4},10^{8},10^{12}$ coincide with the polynomial least squares approximation (PLS) curve and the polynomial limit interpolant (Diag), demonstrating that fixing the condition number (FC) or reusing shape parameters (FS) provides comparable accuracy results.
@@ 3.1.2 $f(x)=\frac{1}{1+25x^{2}}$
Figure 3: Relative $\ell_{2}$ error vs. $N$ for $f(x)=\frac{1}{1+25x^{2}}$ using the $C^{2}(\mathbb{R}^{3})$ Wendland kernel. (Left) Fixed-shape (FS) strategy: Shape parameters $\epsilon$ are tuned on the finest grid and reused. (Right) Fixed condition number (FC) strategy: $\epsilon$ is adjusted per grid to maintain fixed condition numbers ($K_{t}=10^{12},10^{8},10^{4}$).
Figure 3 shows that for both FS and FC the relative $\ell_{2}$ error decays at a root-exponential rate for all kernel smoothness orders and reaches near machine precision. All curves for $K_{t}=10^{4},10^{8},10^{12}$ coincide with the polynomial least squares approximation (PLS) curve and polynomial limit interpolant (Diag), demonstrating that fixing the condition number (FC) or reusing shape parameters (FS) provides comparable accuracy results. Clearly, the unified interpolant behaves purely as a polynomial least squares approximant that happens to interpolate.

### v3

@@ 1 Introduction
$\phi_{m,n}(r)=\begin{cases}\frac{1}{\Gamma(n)2^{n-1}}\int\limits_{r}^{1}s(1-s)^{m}(s^{2}-r^{2})^{n-1}ds&\text{for }0\leq r\leq 1,\\ 0&\text{for }r>1.\end{cases}$ (4)
@@
To situate our approach among nearby methodologies, we briefly contrast it with universal kriging, stabilized global RBF methods, and prior work on compactly-supported RBFs. Our formulation is deterministic and should be distinguished from universal kriging, where a polynomial trend is coupled with a statistically calibrated covariance kernel; see the classic monographs [7, 38]. We instead combine multivariate polynomials with compactly supported positive-definite kernels to obtain a sparse interpolation system that admits efficient numerical linear algebra. In contrast to stabilized global RBF approaches for infinitely smooth kernels (e.g., Gaussian/MQ) that rely on specialized flat-limit algorithms such as RBF-QR or related techniques [19, 17, 9], our focus is on compactly-supported kernels where sparsity, conditioning, and cost are controlled by the support/shape parameter. Finally, within the well-developed literature on compactly-supported RBFs (see, e.g., [6, 40]), our contribution is algorithmic and practical: a unified interpolant with simple, reproducible support selection for the RBF part, polynomial reproduction for some specified space of polynomials, and accompanying numerical linear algebra techniques that are efficient (without a loss in accuracy).
@@
Note: A natural question that arises when reading this work might be “why not use PHS kernels in conjunction with polynomials as global interpolants instead?”. This is indeed a valid choice, but the PHS kernels (used globally) induce a dense kernel Gramian block, which forfeits the sparsity and computational advantages we target here. Further, due their lack of positive-definiteness, they do not admit efficient linear algebra. Even if one worked with a reduced Schur complement system, our approach is asymptotically faster. Our approach also carries over to any positive-definite kernel, but is particularly efficient for compactly-supported ones. Finally, the storage requirements of the unified interpolant presented in this work are significantly lower than those of PHS kernels with polynomials. We do believe the PHS kernels with polynomials are still an appropriate choice for local interpolation and the generation of RBF-based finite difference (RBF-FD) weights.
@@ 2 Unified interpolation with sparse kernels and polynomials
In the following discussion, let $f:\mathbb{R}^{d}\to\mathbb{R}$ be some target function, and let $X=\{\bm{x}_{k}\}_{k=1}^{N}\subset\mathbb{R}^{d}$ be the set of data sites for interpolation where samples of $f$ are provided. Let $\bm{y}=\left.f\right|_{X}$, i.e., $\bm{y}$ is the vector formed by evaluating $f$ at the $N$ data sites $\bm{x}_{k}$. Further, let $q$ be the separation distance in $X$, and $w$ be the largest pairwise distance. Finally, let $r=1/\varepsilon$ be the support of the Wendland kernel. We are primarily interested in two regimes: (1) $r<q$, where the polynomials dominate the approximation, and (2) $q<r<w$, where both kernels and polynomials contribute to the approximation, but the kernels still produce sparse Gramians. Both these regimes admit fast and specialized numerical linear algebra, which carries over also to the case where $r=w$; we do present some experiments for this latter regime in Section 3.
Note: In $\mathbb{R}^{d}$, we use tensor-products of univariate Legendre polynomials restricted to a total degree index set so that they form a basis for the space of polynomials of total degree $\ell$ in dimension $d$. These are evaluated using the usual three-term recurrence relations. Though this is not necessary for the experiments in this paper (which are all on the unit interval or the unit ball), we rescale all points to the unit interval $1,1]^{d}$ so that we can use the standard definitions of Legendre polynomials. We only do this rescaling for the polynomial part of the approximation; the kernels are evaluated on the original point set. We also note that any basis for polynomials can be used instead (even monomials), but we found that tensor products of orthogonal polynomials were easier to work with from a software engineering standpoint.
@@ 2.1.1 Interpolation
When the support $r$ of any Wendland kernel is made smaller than $q$, the shifts of the Wendland kernel each take on the value of $1$ at their center and 0 elsewhere, resembling smooth bump functions. This diagonalizes the kernel Gramian to the $N\times N$ identity matrix $I$; we alternatively refer to the polynomial limit as the diagonal limit for this reason. Thus, the interpolation constraints (2)–(3) enforced at the set of nodes $X$ lead to the following linear system:
@@ 2.1.2 Linear Algebra
    Compute $P=QR$, the (reduced) QR decomposition of the polynomial least-squares matrix $P$ $O(NM^{2})$].
@@
    Solve the Schur complement system for the polynomial coefficients $\mathbf{d}$ as $\mathbf{d}=R^{-1}Q^{T}\bm{y}$ $O(N^{2})+O(NM)$].
@@ 2.2.2 Linear Algebra
    Compute $A=LL^{T}$, the Cholesky decomposition of the RBF matrix AA $O(N^{3})]$.
@@
    Compute the matrix $B=L^{-1}P$ and the vector $\mathbf{g}=L^{-1}\bm{y}$ $O(MN^{2})+O(N^{2})$].
@@
    Compute $B=\tilde{Q}\tilde{R}$, the QR decomposition of $B$ $O(NM^{2})$].
@@
    Solve for the polynomial coefficients $\mathbf{d}$ as $\mathbf{d}=\tilde{R}^{-1}\tilde{Q}^{T}\mathbf{g}$ $O(NM)+O(M^{2})$].
@@
    Solve for the RBF coefficients $\bm{c}$ as $\bm{c}=L^{-T}\left(\mathbf{g}-B\mathbf{d}\right)$ $O(N^{2})$].
@@ 3 Numerical Results
We now present results for interpolation of target functions on Euclidean domains for $d=1,2,3$ and for manifolds $\mathbb{M}\subset\mathbb{R}^{3}$ of co-dimension 1. For $d=1$, the domain is the unit interval [1,1][-1,1] and we use Chebyshev extrema (Chebyshev–Lobatto nodes) as our interpolation data sites (unless otherwise mentioned). For $d=2$, the domain is the unit disk, and for $d=3$ the unit ball. For both $d=2$ and $d=3$, we use boundary-clustered Poisson disk samples generated using the node generation algorithms from [31] as the interpolation data sites, with the points near the boundary being at $0.75h$ instead of $h$ (where $h$ is the fill distance); two representative node sets are shown in Figure 1. These points are clearly non-uniform. The manifolds are point cloud discretizations of the unit sphere and a torus; we present more details in Section 3.4.
Shape parameter $\epsilon$.
Recall that $A(\varepsilon)\in\mathbb{R}^{N\times N}$ is the kernel (Gramian) matrix built with the Wendland kernel at shape parameter $\varepsilon$ on a given node set; we make this dependence on $\varepsilon$ explicit. We write $\kappa(A)$ for the matrix condition number, and $K_{t}\in\{10^{4},10^{8},10^{12}\}$ for a chosen target condition number. We experimented with two different strategies to control the kernel shape parameter $\varepsilon$:
@@
In order to solve for a shape parameter that meets a target condition number, we use the approach outlined in [37]: we rootfind on $\Phi(\epsilon)=\log 10\left(cond(A(\epsilon)\right))-\log 10(K_{t})$. While other approaches for directly selecting $\epsilon$ based directly on internodal distances may be more applicable in practice, this approach allowed us to automate shape parameter calculations without much trial and error.
Polynomial degree $\ell$.
@@
In all experiments, we report the relative $\ell_{2}$ error as
$e_{\ell_{2}}=\frac{\left\|\,s\!\big|_{X_{e}}-f\!\big|_{X_{e}}\right\|_{2}}{\left\|\,f\!\big|_{X_{e}}\right\|_{2}}.$
For $d=1$, we used $2^{14}$ equispaced points on [1,1][-1,1] as the set $X_{e}$. For $d=2,3$, we used densely sampled, quasi-uniform Poisson disk samples with $N_{e}=21748$ and $N_{e}=27987$ respectively, again computed using the node generator from [31]. On manifolds, we used different node generators based on the manifold itself, but set the number of evaluation points to $N_{e}=15000$.
PHS+poly.
We also compare the globally-supported polyharmonic splines with polynomial augmentation (“PHS+poly”) to the unified interpolant and to pure polynomials. For PHS, we take $\phi(r)=r^{2m+1}$ with $p=2m+1\in\{5,7,9\}$ (i.e., $m\in\{2,3,4\}$) and enforce the standard unisolvency condition $\ell\geq m$ for the total-degree polynomial space. The coefficients are obtained from the dense saddle point system
$\begin{bmatrix}K&P\\ P^{\top}&0\end{bmatrix}\begin{bmatrix}c\\ d\end{bmatrix}=\begin{bmatrix}y\\ 0\end{bmatrix}$,
where $K_{ij}=\phi(\|x_{i}-x_{j}\|)$ and $P_{ij}=p_{j}(x_{i})$. In contrast, the unified approach couples a sparse Wendland kernel block with the same polynomial space and is solved by the factorization strategy of Sec. 2.2.2.
To avoid fragmenting the presentation, we integrate PHS+poly into all accuracy and timing plots as a single PHS curve. We select $p=9$ ($m=4$) because (i) across our sweeps $p=9$ delivered the lowest PHS errors and therefore constitutes a best–case PHS baseline under the common degree scaling $\ell=\lfloor C(d)\,N^{1/d}\rfloor$, and (ii) changing $p$ does not alter the qualitative timing order (PHS remains dense and slower in assembly/solve). In 3D, when $p=9$, the KKT system exceeded available memory on the two finest grids; in these cases, we either plot the available $p=9$ levels (truncated abscissa) or substitute $p=7$ solely on those panels to keep abscissae aligned; captions note this explicitly.
Throughout, FS uses the shape parameter fixed at $\epsilon=10$. For each target and dimension we report (i) relative $\ell_{2}$ error on the dense evaluation set, (ii) assembly-and-solve time (linear-solve phase only; formation of kernel/Vandermonde blocks is excluded for all methods), and (iii) evaluation time. Legends label the PHS curve as “PHS+poly (best case, p=9p=9)” and indicate any 3D exceptions.
PHS+poly on manifolds.
On co-dimension–one manifolds $\mathbb{M}\subset\mathbb{R}^{3}$, we also evaluate the globally supported PHS + poly interpolant using the same ambient total-degree polynomial space $\mathcal{P}_{\ell}(\mathbb{R}^{3})$ as in the unified interpolant. However, we now solve the saddle point system in the least-squares sense with a small pseudoinverse tolerance ($10^{-11}$) to account for the rank deficiency of $P$. We swept $m\in\{1,3,5,7,9\}$ and found that on manifolds the lowest errors were consistently achieved at $m=1$ for our degree scaling $\ell=\lfloor C(d)\,N^{1/d}\rfloor$ with $C(\mathbb{M})=1$. Note: this differs from Euclidean tests where we reported a best-case PHS baseline with $p=9$; the manifold geometry and the polynomial scaling lead to a different optimal $p$.
Implementation note. We also implemented a Schur–complement variant based on a row-pivoted LU factorization of the PHS kernel block to eliminate the polynomial coefficients; however, across our tests this approach produced larger interpolation errors than the saddle point solve, so all reported PHS+poly results use the saddle point formulation.
3.1 Univariate functions on [1,1][-1,1]
@@
We investigate the performance of the hybrid RBF-polynomial interpolation with Wendland kernels of smoothness $C^{2}(\mathbb{R}^{3})$, $C^{4}(\mathbb{R}^{3})$, and $C^{6}(\mathbb{R}^{3})$ for the piecewise-smooth function $f(x)=|x|$). Since this target lacks smoothness, we expect slow convergence rates for all our methods. The increasing polynomial degrees for this case are as follows: $4$, 88, $16$, $32$, $64$, $128$, $256$.
Figure 2: Relative $\ell_{2}$ error vs. $N^{1/d}$ (here, NN) for $f(x)=|x|$ using the $C^{2}(\mathbb{R}^{3})$ Wendland kernel. (Left) Fixed-shape (FS) strategy: Shape parameters $\epsilon$ are tuned on the finest grid and reused. (Right) Fixed condition number (FC) strategy: $\epsilon$ is adjusted per grid to maintain fixed condition numbers ($K_{t}=10^{12},10^{8},10^{4}$).
Figure 2 shows that for both FS and FC the relative $\ell_{2}$ error decays algebraically for the $C^{2}(\mathbb{R}^{3})$ kernel; we observed similar results for kernels of higher smoothness as well. This is consistent with the lack of smoothness of the target function. Note that all curves for $K_{t}=10^{4},10^{8},10^{12}$ coincide with the polynomial least squares approximation (PLS) curve and the polynomial limit interpolant (Diag), demonstrating that fixing the condition number (FC) or reusing shape parameters (FS) provides comparable accuracy results. We also note that PHS+poly curve behaves much worse on both figures.
Figure 3: Computational costs vs. NN for $f(x)=|x|$ using the $C^{2}(\mathbb{R}^{3})$ Wendland kernel. (Top row) Fixed-shape (FS) strategy: Shape parameters $\epsilon$ are tuned on the finest grid and reused. (Bottom row) Fixed condition number (FC) strategy: $\epsilon$ is adjusted per grid to maintain fixed condition numbers ($K_{t}=10^{12},10^{8},10^{4}$). Left: assembly and solve time; Right: evaluation time.
@@ 3.1.2 $f(x)=\frac{1}{1+25x^{2}}$
Figure 4: Relative $\ell_{2}$ error vs. NN for $f(x)=\frac{1}{1+25x^{2}}$ using the $C^{2}(\mathbb{R}^{3})$ Wendland kernel. (Left) Fixed-shape (FS) strategy: Shape parameters $\epsilon$ are tuned on the finest grid and reused. (Right) Fixed condition number (FC) strategy: $\epsilon$ is adjusted per grid to maintain fixed condition numbers ($K_{t}=10^{12},10^{8},10^{4}$).
Figure 4 shows that for both FS and FC the relative $\ell_{2}$ error decays at a root-exponential rate for all kernel smoothness orders and reaches near machine precision. All curves for $K_{t}=10^{4},10^{8},10^{12}$ coincide with the polynomial least squares approximation (PLS) curve and polynomial limit interpolant (Diag), demonstrating that fixing the condition number (FC) or reusing shape parameters (FS) provides comparable accuracy results. Clearly, the unified interpolant behaves purely as a polynomial least squares approximant that happens to interpolate. Similarly to Figure 2, Figure 4 shows PHS+poly curve behaving worse than the rest of the experiments.
Figure 5: Computational costs vs. NN for $f(x)=\frac{1}{1+25x^{2}}$ using the $C^{2}(\mathbb{R}^{3})$ Wendland kernel. (Top row) Fixed-shape (FS) strategy: Shape parameters $\epsilon$ are tuned on the finest grid and reused. (Bottom row) Fixed condition number (FC) strategy: $\epsilon$ is adjusted per grid to maintain fixed condition numbers ($K_{t}=10^{12},10^{8},10^{4}$). Left: assembly and solve time; Right: evaluation time.
3.1.3 Timings.
We first report computational cost as a function of $N$ on the same boundary–clustered interval node sets (Chebyshev interior points with endpoints). Figure 3 reports wall–clock cost versus $N$ for $f(x)=|x|$. Two clear patterns appear. First, evaluation time is essentially identical across PLS, Diag, and FS, with gently increasing curves as $N$ grows; once coefficients are computed, all three methods evaluate the same hybrid basis, so no systematic advantage is expected. In contrast, PHS+poly evaluates more slowly due to dense global kernel evaluation and therefore sits above the clustered PLS/Diag/FS bundle in the right panels. Second, assembly and solve separates the methods: PLS and Diag track one another closely and remain the cheapest across all $N$, while FS is consistently more expensive with a visibly steeper slope due to neighbor searches and CSRBF block assembly, and the denser local stencils implied by the fixed-shape choice tuned on the finest grid. PHS+poly lies between these groups: its dense saddle point system makes it more expensive than PLS/Diag but still below the FS curve (left panels). In the FC panels, the same ordering holds and the family of curves is monotone in the prescribed condition number: higher $K_{t}$ (e.g., $10^{12}$) generates larger supports and noticeably higher assembly cost than $K_{t}=10^{8}$ and $10^{4}$, whereas evaluation time remains clustered for all $K_{t}$. Small $N$ transients cause mild crossings near the left edge, but the asymptotic ordering is stable.
Figure 5 shows the same timing breakdown for the analytic Runge function. The qualitative story matches the nonsmooth case. Evaluation curves for PLS, Diag, and FS lie on top of each other for all $N$, reflecting identical evaluation kernels once coefficients are known, while PHS+poly again evaluates more slowly and forms the upper curve in the right panels. Assembly and solve again favors PLS and Diag, with FS incurring the highest cost; PHS+poly remains intermediate between PLS/Diag and FS, consistent with previous results. Under FC, assembly time increases systematically with the target condition number ($K_{t}=10^{12}>10^{8}>10^{4}$), while evaluation time is largely insensitive to $K_{t}$. Across both FS and FC, PLS and Diag remain virtually indistinguishable, FS exhibits a consistent assembly premium that widens with $N$, PHS+poly remains strictly above PLS/Diag yet below FS, and all methods display smooth, near-linear growth on the log-log scale.